OpenStudy (anonymous):

Medal and fan. What are the real or imaginary solutions to this equation x^4-52x^2+576=0

OpenStudy (anonymous):

\[x^4-52x^2+576=0\] does it factor?

OpenStudy (anonymous):

Again (i'm bad at factoring) I have to find something that adds to -52 and multiplies to 576?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

try \(16\) and \(36\)

OpenStudy (anonymous):

So -16 and -36, How you came up with that that quickly I will never know.

OpenStudy (anonymous):

i cheated i suck at factoring but we can do this without factoring if you like might even be easier

OpenStudy (anonymous):

Eh, I should probably learn to factor if that's alright?

OpenStudy (anonymous):

Or at least practice it.

OpenStudy (anonymous):

well you should really learn how to solve a quadratic equation without factoring most of them do not factor using whole numbers in any case

OpenStudy (anonymous):

Using the quadratic formula yeah?

OpenStudy (anonymous):

and how you would know to come up with 16 and 36 is anyone’s guess i would not use the quadratic formula for this one i would complete the square, much easier here

OpenStudy (anonymous):

Okay, there's another way lol. Teach me how to complete the square? /: I'll be your friend forever.

OpenStudy (anonymous):

ok lets start with an easy one then we do this one ok?

OpenStudy (anonymous):

Sounds good.

OpenStudy (anonymous):

lets start with \[x^2+6x-5=0\] add \(5\) to both sides, what do you get?

OpenStudy (anonymous):

Just double checking, It will give me both answers right? I know this calls for two.

OpenStudy (anonymous):

x^2+6x=5

OpenStudy (anonymous):

yeah we will finish this in a moment ok now you have \[x^2+6x=5\] what is half of \(6\) ?

OpenStudy (anonymous):

3.

OpenStudy (anonymous):

right what is \(3^2\)?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

ok good so next step is to turn \[x^2+6x=5\] in to \[(x+3)^2=5+9\\ (x+3)^2=14\]

OpenStudy (anonymous):

Yay. Good luck me.

OpenStudy (anonymous):

a word of explanation \[(x+3)^2=x^2+6x+9\] so by adding 9 to both sides you changed \[x^2+6x=5\] to \[(x+3)^2=14\]

OpenStudy (anonymous):

that makes \[x+3=\pm\sqrt{14}\] and so \[x=-3\pm\sqrt{14}\]

OpenStudy (anonymous):

notice you could never in a million years factor \[x^2+6x-5=0\]

OpenStudy (anonymous):

Okay so. Now I need to know the sqrt of 14? And solve

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

yes that is right when you take the square root, don't forge the \(\pm\)

OpenStudy (anonymous):

So 3 +- 3.7

OpenStudy (anonymous):

So 6.7 and -.7?

OpenStudy (anonymous):

leave it with the radical \[3\pm\sqrt{14}\] looks nicer forget the decimals

OpenStudy (anonymous):

but yeah you are right

OpenStudy (anonymous):

Okay, But on the next I have to solve it out.

OpenStudy (anonymous):

shall we finish this problem? the one you posted?

OpenStudy (anonymous):

Yes, please.

OpenStudy (anonymous):

well that was wrong in any case

OpenStudy (anonymous):

do you want to believe it factors as \[(x^2-36)(x^2-16)=0\] or do you want to complete the square again?

OpenStudy (anonymous):

Lol! I was about to feel incredibly smart for finding a typo! It factors as (x-16)(x-36) yeah?

OpenStudy (anonymous):

Can we solve this way?

OpenStudy (anonymous):

yeah a typo but you are probably incredibly smart in any case

OpenStudy (skullpatrol):

psst...difference of two squares ;-)

OpenStudy (anonymous):

Wellllllllll, thanks for that (:

OpenStudy (anonymous):

the factor way is easy if we know if factors both \(x^2-36\) and \(x^2-16\) are the difference of two squares

OpenStudy (anonymous):

so factor all the way down and you get the zeros, more or less in your head

OpenStudy (anonymous):

The difference of two squares? I'm confused. Don't lose me just yet.

OpenStudy (anonymous):

\[a^2-b^2=(a+b)(a-b)\] \[x^2-36=(x+6)(x-6)\]

OpenStudy (anonymous):

similar for \(x^2-16\)

OpenStudy (skullpatrol):

$$\Huge a^2-b^2=(a+b)(a-b)$$

OpenStudy (anonymous):

So that's the equation to find the difference of two squares?

OpenStudy (anonymous):

yeah i always factors that way of course if you had \[x^2-16=0\] and you were a thinking human being you would write \[x^2=16\\x=4,x=-4\]

OpenStudy (anonymous):

Oh wait! okay I see that. So then the equation for x^2-16 would be x^2-16=(x+4)(x-4)

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

Ha! I'm a thinking human then. So are those my answers? Though How do I know which two?

OpenStudy (anonymous):

not two , four!

OpenStudy (anonymous):

So they're all four my answer?

OpenStudy (anonymous):

Okay. I see that. Can I get your help on one more (it's similiar)? /:

OpenStudy (anonymous):

\[x^4-52x^2+576=0\] a polynomial of degree 4 this one has 4 real zeros once you have it factored all the way down to its bare bones it is \[(x+6)(x-6)(x+4)(x-4)=0\]

OpenStudy (anonymous):

as you can see this problem has been cooked to within an inch of its life to give such nice answers mostly they don't work out so nicely

OpenStudy (anonymous):

sure go ahead, we can do another one

OpenStudy (anonymous):

Okay. This one says the same, to find the real and imaginary answers of x^3=216

OpenStudy (skullpatrol):

$$\Huge ab=0 \iff a=o \text{ or } b=0.$$

OpenStudy (anonymous):

this one is different, you want to start with \[x^2-216\] and factor as the difference of two cubes

OpenStudy (anonymous):

did you mean x^3? Or..

OpenStudy (anonymous):

of course in order to do that you have to a) recognize that \(216=6^3\) and that b) \(a^3-b^3=(a-b)(a^2+ab+b^2)\)

OpenStudy (anonymous):

Alright, this is tough. Walk me through it and I'll do the math. Just point the way.

OpenStudy (anonymous):

yeah i meant \[\huge x^{\color{red}3}-216=0\]

OpenStudy (anonymous):

\[x^3-6^3=0\] is how you visualize it

OpenStudy (anonymous):

then use \[a^3-b^3=(a-b)(a^2+ab+b^2)\] with \[a=x,b=6\]

OpenStudy (anonymous):

Okay. So. \[x^3-6^3=(x-6)(x^2+6x+6^2) \]

OpenStudy (anonymous):

yes aka \[(x-6)(x^2+6x+36)=0\]

OpenStudy (anonymous):

Okay. You're very fast at typing equations, if only there was a medal. (;

OpenStudy (anonymous):

now one answer is obvious \[x-6=0\\ x=6\] the other two complex zeros you get are from solving \[x^2+6x+36=0\] the solutions are not real you want to find them?

OpenStudy (anonymous):

I can obviously try lol. Lets see..

OpenStudy (anonymous):

\[x^2+6x=-3\]

OpenStudy (anonymous):

-36**

OpenStudy (anonymous):

\[x^2+x=-6\]

OpenStudy (anonymous):

Am I doing this right. Because I really don't think I am.

OpenStudy (anonymous):

ok good then \[x^2+6x=-36\]is the first step next step what is half of 6?

OpenStudy (anonymous):

3.

OpenStudy (anonymous):

and \(3^2\)>

OpenStudy (anonymous):

9?

OpenStudy (anonymous):

no, not 9?, but 9! so we go from \[x^2+6x=-36\] to \[(x+3)^2=-36+9\\ (x+3)^2=-27\]

OpenStudy (anonymous):

like i said, the answers are complex because now when you take the square root you will be taking the square root of a negative number

OpenStudy (anonymous):

Okay. Yeah, imaginary numbers.

OpenStudy (anonymous):

you get \[x+3=\sqrt{-27}\] you know how to write that as a complex number?

OpenStudy (anonymous):

actually you get \[x+3=\pm\sqrt{-27}\]

OpenStudy (anonymous):

Uhh lol would it be 3+ something I and 3-something i

OpenStudy (anonymous):

actually since \(27=9\times 3\) it is \[\sqrt{-27}=3\sqrt{3}i\]

OpenStudy (anonymous):

making your final answer \[-3\pm3\sqrt{3}i\]

OpenStudy (anonymous):

Alright. I think I understand heh. Thank you so much!

OpenStudy (anonymous):

yw